A rectangle has length 6cm greater than its width. Find its width given that its area is 91cm^2
Answer by bobdabuildr(27)

Width = x
Length = 6+x
Area = length X width
91=x(x+6)
91=x^2 + 6x
0= x^2 +6x-91
0=(x+13)(x-7)
x=-13 or x=7
Since distance can not be negative the x=-13 is an erroneous answer
Check
91=7(6+7)
91=91